# AgentDR: Dynamic Recommendation with Implicit Item-Item Relations via LLM-based Agents

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Retrieval models and ranking
@@ 1. Introduction
Recommender systems provide personalized suggestions to assist users in information discovery and decision-making (Wang et al., 2021; Zhang et al., 2019; Wu et al., 2023b). However, traditional ID-based recommender systems (He et al., 2020; Kang and McAuley, 2018; Chen et al., 2020b) solely based on user-item interactions are constrained by overlooking the rich semantic information contained in textual context. Even when LLMs are employed to encode textual features into embeddings (Wei et al., 2024; Ren et al., 2024), these fixed-length representations inherently struggle to preserve the full semantic richness of textual information for recommendation (Li et al., 2025; Wu et al., 2024b). There are several attempts to harness the commonsense reasoning and knowledge utilization capabilities of LLMs by directly applying them as recommender systems, reformulating recommendation tasks as language modeling problems (Geng et al., 2022; Hou et al., 2024; Cui et al., 2022). However, an inherent gap remains because LLMs are optimized for next-token prediction in natural language rather than learning collaborative and sequential patterns from interactions, limiting their efficiency in user behavior modeling.
Figure 1. Two major drawbacks in directly deploying LLMs for recommendation: LLMs can hallucinate items not existing in the actual product catalog, and token limits of LLMs make them unsuitable for ranking items at scale. AgentDR addresses both limitations by delegating full-ranking tasks to recommendation tools.
Given the limitations of using LLMs solely as feature encoders or standalone recommenders, recent studies have investigated leveraging LLMs as interactive agents for recommendation (Zhang et al., 2024b, a; Wang et al., 2025b). Rather than relying only on direct prompting with user interaction histories, these agent-based approaches integrate memory modules to preserve long-term context and better harness collaborative signals throughout the recommendation process. While LLM-based agents alleviate the representational constraints of static numerical embeddings and enable adaptive modeling of user behavior over time, they also inherit key limitations from the generative nature of LLMs. As illustrated in Fig. 1, two primary drawbacks arise in previous agent-based works when directly applying LLMs for recommendation: (i) the tendency to hallucinate non-existent items that are not part of the product catalog, and (ii) the high inference cost and token length constraints, which make them impractical for large-scale item ranking. Moreover, longer input and output sequences further increase the risk of hallucination due to attention bottlenecks and heightened generation uncertainty (Huang et al., 2025c).
@@
Besides their limitations in full-ranking, previous LLM-based agents primarily focus on modeling user preferences from explicit user-item interactions, which already can be effectively modeled by traditional embedding-based recommendation methods (Mao et al., 2021; Chen et al., 2020b; He et al., 2020). On the contrary, traditional models often overlook higher-level semantic relations between items that go beyond co-occurrence patterns. To address this gap, the generative capabilities and extensive world knowledge of LLMs make them particularly well-suited for reasoning over implicit relationships, such as complementarity and substitution relations. Substitutes help identify alternative options when a preferred item is unavailable or when the user seeks variety, while complements reveal co-purchase patterns that often indicate planned or contextual consumption behaviors. For example, a user purchasing a digital single-lens reflex camera often requires a tripod as a complement, whereas a user who buys tiramisu is more likely to choose cheesecake as substitute over burgers when the original item is out of stock. Although such relationships can greatly enhance recommendation quality (Zhang et al., 2022; Wu et al., 2023a), the ground-truth labels of them are typically absent from most datasets (Ni et al., 2019; Bonab et al., 2021a). Furthermore, annotating item pairs with accurate relational labels often requires extensive world knowledge and nuanced understanding, making manual annotation both expensive and time-consuming. This gap highlights the potential of LLMs as an alternative source of relational knowledge. By leveraging their extensive pretraining on diverse textual data, LLMs can capture subtle functional relationships between items, enabling automatic discovery of substitutes and complements that would otherwise require labor-intensive human annotation.
@@ 2. Related Works
Method Capability Full Ranking
Evaluation Text Feature
Handling Generative
Reasoning
@@
LLM-Rec Agent (Huang et al., 2025b; Wang et al., 2024; Zhang et al., 2024b, a; Wang et al., 2025b; Xu et al., 2025)
@@
LLM-based agents have recently demonstrated strong capabilities in using memory and reflection to learn from past experiences and simulate human-like behaviors (Gao et al., 2025; Zhong et al., 2024; Shinn et al., 2023; Zhao et al., 2024; Shi et al., 2024; Li et al., 2024). In the context of recommendation, LLM-based agents have been used to investigate social phenomena such as information cocoons and filter bubble effect by simulating user behaviors (Wang et al., 2025b; Zhang et al., 2024a). Given that agent-based frameworks can offer deeper insights into user behavior by leveraging the reasoning capabilities of LLMs, there is a growing interest in applying LLM-based agents to enhance recommendation performance (Wang et al., 2024; Huang et al., 2025b; Zhang et al., 2024b; Xu et al., 2025). Nonetheless, since hallucination (Yu et al., 2024; Shen et al., 2025) and token limitation remain two challenges when applying these agents to rank all items in the datasets for each user, most prior works (Zhang et al., 2024b; Wang et al., 2024; Huang et al., 2025b; Zhang et al., 2024a; Wang et al., 2025b; Xu et al., 2025) lack of full-ranking ability. The batch-ranking task in these works is fundamentally misaligned with most real-world scenarios, where recommender systems are required to rank relevant items from massive catalogs. And these works focus on predicting user behaviors based on user-item interactions, a signal that traditional recommendation models (He et al., 2020; Kang and McAuley, 2018; Mao et al., 2021) are already well-equipped to capture and exploit. Different from these works, our framework combines the strengths of LLMs and traditional recommenders to achieve personalized full-ranking. Related works other than LLM-based agent frameworks for recommendation are discussed in Appendix  A.4. Comparison between different approaches is summarized in Table 1.
@@ 3. Proposed Framework: AgentDR
Figure 2. The framework of AgentDR. Each user agent is equipped with two memory modules: RecTool memory to store tool suitability, and intent memory to track user intent. Ranking results from recommendation tools are aggregated based on RecTool memory. The aggregated result is further refined by dual S&C and general ranking modules.
@@ 3.2.1. User Profile Summarization
(1) $y_{\text{prof}}=\text{LLM}_{\text{prof}}(\text{desc}(\mathbf{s}))$
@@ 3.3. Substitute and Complement Generation
(3) $y_{\text{gen}}^{com}$ $=\text{LLM}_{\text{gen}}^{com}(\text{desc}(\mathbf{s}_{[-(k+c):-k]}))$
@@ 3.4.1. LLM-based Tool Comparison
(4) $\mathbf{z}_{\text{cpr}}=\text{LLM}_{\text{cpr}}(\{\text{desc}(f_{\text{top}}(\hat{\mathbf{r}}_{t},k_{\text{cpr}}))|t\in\mathcal{T}\},\text{desc}(\mathbf{s}_{[-k:]}))$
@@
(5) $m_{t}\leftarrow m_{t}+\alpha\cdot z_{\text{cpr}}^{t}$
@@ 3.4.2. Ranking Comparison and Aggregation
(6) $m_{t}\leftarrow m_{t}+\beta\sum_{\{r^{j}_{t}|i_{j}\in\mathbf{s}_{[-k:]}\}}\frac{(r^{j}_{t})^{-1}}{\sum_{t^{\prime}\in\mathcal{T}}(r^{j}_{t^{\prime}})^{-1}}$
@@
(7) $\hat{\mathbf{r}}=\sum_{t\in\mathcal{T}}m_{t}\hat{\mathbf{r}}_{t}=m_{G}\hat{\mathbf{r}}_{G}+m_{S}\hat{\mathbf{r}}_{S}+m_{M}\hat{\mathbf{r}}_{M}$
@@ 3.5.1. Substitution and Complementarity
(8) $\mathbf{z}_{\text{dcm}}=\text{LLM}_{\text{dcm}}(y_{\text{gen}}^{sub},y_{\text{gen}}^{com},\text{desc}(\mathbf{s}_{[-k:]}))$
@@ 3.5.3. General Reranking
(10) $y_{\text{rank}}^{reg}=\text{LLM}^{reg}_{\text{rank}}(f_{\text{top}}(\hat{\mathbf{r}},k^{\prime}),\text{desc}(f_{\text{top}}(\hat{\mathbf{r}},k^{\prime})),y_{\text{prof}})$
@@ 3.6. Dual S&C Reranking
(13) $y_{\text{rank}}^{com}$ $=\text{LLM}_{\text{rank}}(f_{\text{top}}(\hat{\mathbf{r}},k^{\prime}),\text{desc}(f_{\text{top}}(\hat{\mathbf{r}},k^{\prime})),y_{\text{gen}}^{com})$
@@
(14) $\mathbf{\hat{r}}^{dual}=f_{\text{fuse}}((m_{sub},y_{\text{rank}}^{sub}),(m_{com},y_{\text{rank}}^{com}))$
@@
(15) $\hat{r}_{i}^{dual}=m_{sub}\cdot\text{index}(i,y_{\text{rank}}^{sub})+m_{com}\cdot\text{index}(i,y_{\text{rank}}^{com})$
@@
(16) $\mathbf{\hat{r}}^{final}=f_{\text{fuse}}((1,y_{\text{rank}}^{dual}),(m_{reg},y_{\text{rank}}^{reg}))$
@@ 4. Experiment
Through our extensive empirical analysis, we aim to address the following research questions. RQ1: How does AgentDR perform in full-ranking recommendation compared to other baselines? RQ2: Does incorporating LLM-based reasoning modules improve the semantic alignment of recommended items with user intent? RQ3: How do different ranking aggregation strategies affect the performance of AgentDR? RQ4: What is the individual contribution of each LLM-based module in AgentDR?
@@ 4.1.1. Datasets
We conduct experiments on three datasets: Instacart [^1] https://www.kaggle.com/yasserh/datasets, Electronics and Sports [^2] https://xmrec.github.io/data/us/. Instacart is a public dataset containing transactional data from the Instacart app, primarily focused on grocery items for basket recommendation. Electronics and Sports are two sub-datasets from public XMarket dataset (Bonab et al., 2021b). Electronics contains user history of purchasing electronics on Amazon. Sports contains user interactions with items under category of sports and outdoors on Amazon. The statistics are summarized in Table 2.
@@ 4.1.2. Tool selection
In this work, three pretrained recommendation models are used as tools within AgentDR:
  1. (1)
    LightGCN (He et al., 2020). A graph-based collaborative filtering model that simplifies traditional graph convolution network by removing feature transformations and nonlinearities.
  2. (2)
    SASRec (Kang and McAuley, 2018). A sequential recommendation model based on Transformer, which captures the user’s dynamic preferences by modeling the dependencies between historical items.
  3. (3)
    SimpleX (Mao et al., 2021). A MF-based model where a margin is applied to filter uninformative negative samples in the loss.
The simplicity of these tools allows us to validate the effectiveness of our framework without relying on complex architectures, aligning with our goal of maintaining a model-agnostic design.
@@ 4.1.3. Baselines
    ENMF (Chen et al., 2020b) introduces mathematical optimization that enable efficient full-data MF without negative sampling.
@@
Both tools and recommendation baselines are pretrained on the first (nk)(n-k) items of each user’s interaction sequence $\mathbf{s}_{[:-k]}$, following the standard training protocol of sequential recommendation for fair performance comparison. Next, following the previous works (Zhang et al., 2024b; Xu et al., 2025), we compare AgentDR with language-based approaches below:
@@
In addition to these language-based baselines, we compare AgentDR against recommendation tools integrated with Retrieval-Augmented Generation (RAG). Specifically, the LLM is prompted to generate the top 20 items from the 50 candidates retrieved by each tool, conditioned on the user’s behavior sequence. The same hallucination filtering mechanism from Sec. 3.6 is applied for fair comparison. These baselines are represented as (9)-(11) RAGG/S/M for LightGCN, SASRec and SimpleX, respectively. Since most prior LLM-agent works (Zhang et al., 2024b; Wang et al., 2024; Huang et al., 2025b; Zhang et al., 2024a; Wang et al., 2025b; Xu et al., 2025) are unable to rank all items in datasets, we exclude them from the comparison table instead of reporting trivially low full-ranking performance.
@@ 4.2. Semantic Vicinity Evaluation: VDCG
Specifically, we prompt the LLM to rate how well each item in the recommended list aligns with the ground-truth item in terms of relevance, usefulness, and user interest. The rating scale ranges from 0 to 9, where 0 indicates an entirely unrelated item, and 9 represents a perfect semantic match between the item descriptions. These ratings are treated as relevance scores to compute the Discounted Cumulative Gain (DCG) (Burges et al., 2005) for the list. For normalization, we divide by a customized ideal DCG, calculated from a synthetic ideal list of scores: $p,p/2,...,p/2^{l}]$, where $p$ is the maximum relevance score minus 1 (i.e., 8), and $l$ is the length of the list. This ideal score distribution reflects the empirical tendency of the LLM to assign higher scores to only a few semantic-related items, following an approximately exponential decay. VDCG thus captures both the semantic vicinity and ranking quality of the predicted list with respect to the user’s true intent.

### v2

@@ Abstract.
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emirates.booktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesisbn: 979-8-4007-2307-0/2026/04doi: 10.1145/3774904.3792304ccs: Information systems Retrieval models and ranking
@@ 1. Introduction
Recommender systems provide personalized suggestions to assist users in information discovery and decision-making (Yang et al., 2023b, 2024b, a; Loveland et al., 2025). However, traditional ID-based recommender systems (He et al., 2020; Kang and McAuley, 2018; Chen et al., 2020a) solely based on user-item interactions are constrained by overlooking the rich semantic information contained in textual context. Even when LLMs are employed to encode textual features into embeddings (Wei et al., 2024; Ren et al., 2024; Zhu et al., 2025), these fixed-length representations inherently struggle to preserve the full semantic richness of textual information for recommendation (Li et al., 2025; Wu et al., 2024b). There are several attempts to harness the commonsense reasoning and knowledge utilization capabilities of LLMs by directly applying them as recommender systems, reformulating recommendation tasks as language modeling problems (Geng et al., 2022; Hou et al., 2024; Cui et al., 2022; He et al., 2025). However, an inherent gap remains because LLMs are optimized for next-token prediction in natural language rather than learning collaborative and sequential patterns from interactions, limiting their efficiency in user behavior modeling.
Figure 1. There are two major drawbacks in directly deploying LLMs for recommendation: hallucination and token limitation. AgentDR addresses them by delegating full-ranking tasks to recommendation tools.
Given the limitations of using LLMs solely as feature encoders or standalone recommenders, recent studies have investigated leveraging LLMs as interactive agents for recommendation (Zhang et al., 2024b, a; Wang et al., 2025a). Rather than relying only on direct prompting with user interaction histories, these agent-based approaches integrate memory modules to preserve long-term context and better harness collaborative signals throughout the recommendation process. While LLM-based agents alleviate the representational constraints of static numerical embeddings and enable adaptive modeling of user behavior over time, they also inherit key limitations from the generative nature of LLMs. As illustrated in Fig. 1, two primary drawbacks arise in previous agent-based works when directly applying LLMs for recommendation: (i) the tendency to hallucinate non-existent items that are not part of the product catalog, and (ii) the high inference cost and token length constraints, which make them impractical for large-scale item ranking. Moreover, longer input and output sequences further increase the risk of hallucination due to attention bottlenecks and heightened generation uncertainty (Huang et al., 2025b).
@@
Besides their limitations in full-ranking, previous LLM-based agents primarily focus on modeling user preferences from explicit user-item interactions, which already can be effectively modeled by traditional embedding-based recommendation methods (Mao et al., 2021; Chen et al., 2020a; He et al., 2020). On the contrary, traditional models often overlook higher-level semantic relations between items that go beyond co-occurrence patterns. To address this gap, the generative capabilities and extensive world knowledge of LLMs make them particularly well-suited for reasoning over implicit relationships, such as complementarity and substitution relations. Substitutes help identify alternative options when a preferred item is unavailable or when the user seeks variety, while complements reveal co-purchase patterns that often indicate planned or contextual consumption behaviors. For example, a user purchasing a digital single-lens reflex camera often requires a tripod as a complement, whereas a user who buys tiramisu is more likely to choose cheesecake as substitute over burgers when the original item is out of stock. Although such relationships can greatly enhance recommendation quality (Zhang et al., 2022; Wu et al., 2023; Nerrise et al., 2025), the ground-truth labels of them are typically absent from most datasets (Ni et al., 2019; Bonab et al., 2021b). Furthermore, annotating item pairs with accurate relational labels often requires extensive world knowledge and nuanced understanding, making manual annotation both expensive and time-consuming. This gap highlights the potential of LLMs as an alternative source of relational knowledge. By leveraging their extensive pretraining on diverse textual data, LLMs can capture subtle functional relationships between items, enabling automatic discovery of substitutes and complements that would otherwise require labor-intensive human annotation.
@@ 2. Related Work
Full Ranking Text Feature Generative
Evaluation Handling Reasoning
@@
LLM-Rec Agents (Huang et al., 2025c; Wang et al., 2024; Zhang et al., 2024b, a; Wang et al., 2025a; Xu et al., 2025)
@@
LLM-based agents have recently demonstrated strong capabilities in using memory and reflection to learn from past experiences and simulate human-like behaviors (Gao et al., 2025; Zhong et al., 2024; Shinn et al., 2023; Zhao et al., 2024; Shi et al., 2024; Li et al., 2024). In the context of recommendation, LLM-based agents have been used to investigate social phenomena such as information cocoons and filter bubble effect by simulating user behaviors (Wang et al., 2025a; Zhang et al., 2024a). Given that agent-based frameworks can offer deeper insights into user behavior by leveraging the reasoning capabilities of LLMs, there is a growing interest in applying LLM-based agents to enhance recommendation performance (Wang et al., 2024; Huang et al., 2025c; Zhang et al., 2024b; Xu et al., 2025). Nonetheless, since hallucination (Yu et al., 2024; Shen et al., 2025) and token limitation remain two challenges when applying these agents to rank all items in the datasets for each user, most prior works (Zhang et al., 2024b; Wang et al., 2024; Huang et al., 2025c; Zhang et al., 2024a; Wang et al., 2025a; Xu et al., 2025) lack of full-ranking ability. The batch-ranking task in these works is fundamentally misaligned with most real-world scenarios, where recommender systems are required to rank relevant items from massive catalogs. And these works focus on predicting user behaviors based on user-item interactions, a signal that traditional recommendation models (He et al., 2020; Kang and McAuley, 2018; Mao et al., 2021) are already well-equipped to capture and exploit. Different from these works, our framework combines the strengths of LLMs and traditional recommenders to achieve personalized full-ranking. Related works other than LLM-based agent frameworks for recommendation are discussed in Appendix  A.5. Comparison between different approaches is summarized in Table 1.
@@ 3. Proposed Framework: AgentDR
Figure 2. The framework of AgentDR: Each user agent is equipped with two memory modules: RecTool memory to store tool suitability, and intent memory to track user intent. Ranking results from recommendation tools are aggregated based on RecTool memory. The aggregated result is further refined by the dual S&C and general ranking modules.
@@ 3.2.1. User Profile Summarization
(1) $y_{\text{prof}}=\text{LLM}_{\text{prof}}(\text{desc}(\mathbf{s})),$
@@ 3.3. Substitute and Complement Generation
(3) $y_{\text{gen}}^{com}$ $=\text{LLM}_{\text{gen}}^{com}(\text{desc}(\mathbf{s}_{[-(k+c):-k]})),$
@@ 3.4.1. LLM-based Tool Comparison
(4) $\mathbf{z}_{\text{cpr}}=\text{LLM}_{\text{cpr}}(\{\text{desc}(f_{\text{top}}(\hat{\mathbf{r}}_{t},k_{\text{cpr}}))|t\in\mathcal{T}\},\text{desc}(\mathbf{s}_{[-k:]})),$
@@
(5) $m_{t}\leftarrow m_{t}+\alpha\cdot z_{\text{cpr}}^{t},$
@@ 3.4.2. Ranking Comparison and Aggregation
(6) $m_{t}\leftarrow m_{t}+\beta\sum_{\{r^{j}_{t}|i_{j}\in\mathbf{s}_{[-k:]}\}}\frac{(r^{j}_{t})^{-1}}{\sum_{t^{\prime}\in\mathcal{T}}(r^{j}_{t^{\prime}})^{-1}},$
@@
(7) $\hat{\mathbf{r}}=\sum_{t\in\mathcal{T}}m_{t}\hat{\mathbf{r}}_{t}=m_{G}\hat{\mathbf{r}}_{G}+m_{S}\hat{\mathbf{r}}_{S}+m_{M}\hat{\mathbf{r}}_{M},$
@@ 3.5.1. Substitution and Complementarity
(8) $\mathbf{z}_{\text{dcm}}=\text{LLM}_{\text{dcm}}(y_{\text{gen}}^{sub},y_{\text{gen}}^{com},\text{desc}(\mathbf{s}_{[-k:]})),$
@@ 3.5.3. General Reranking
(10) $y_{\text{rank}}^{reg}=\text{LLM}^{reg}_{\text{rank}}(f_{\text{top}}(\hat{\mathbf{r}},k^{\prime}),\text{desc}(f_{\text{top}}(\hat{\mathbf{r}},k^{\prime})),y_{\text{prof}}),$
@@ 3.6. Dual S&C Reranking
(13) $y_{\text{rank}}^{com}$ $=\text{LLM}_{\text{rank}}(f_{\text{top}}(\hat{\mathbf{r}},k^{\prime}),\text{desc}(f_{\text{top}}(\hat{\mathbf{r}},k^{\prime})),y_{\text{gen}}^{com}),$
@@
(14) $\mathbf{\hat{r}}^{dual}=f_{\text{fuse}}((m_{sub},y_{\text{rank}}^{sub}),(m_{com},y_{\text{rank}}^{com})),$
@@
(15) $\hat{r}_{i}^{dual}=m_{sub}\cdot\text{index}(i,y_{\text{rank}}^{sub})+m_{com}\cdot\text{index}(i,y_{\text{rank}}^{com}),$
@@
(16) $\mathbf{\hat{r}}^{final}=f_{\text{fuse}}((1,y_{\text{rank}}^{dual}),(m_{reg},y_{\text{rank}}^{reg})),$
@@ 4. Experiment
Through our extensive empirical analysis, we aim to address the following research questions. (RQ1) How does AgentDR perform in full-ranking recommendation compared to other baselines? (RQ2) Does incorporating LLM-based reasoning modules improve the semantic alignment of recommended items with user intent? (RQ3) How do different ranking aggregation strategies affect the performance of AgentDR? (RQ4) What is the individual contribution of each LLM-based module in AgentDR?
@@ 4.1.1. Datasets
We conduct experiments on three datasets: Instacart (Kaggle, 2022), Electronics and Sports (Bonab et al., 2021a). Electronics and Sports are two sub-datasets from public XMarket dataset (Bonab et al., 2021c). The statistics are summarized in Table 2.
@@ 4.1.2. Tool selection
In this work, three pretrained recommendation models are used as tools within AgentDR: (1) LightGCN (He et al., 2020), (2) SASRec (Kang and McAuley, 2018), and (3) SimpleX (Mao et al., 2021). The simplicity of these tools allows us to validate the effectiveness of our framework without relying on complex architectures, aligning with our goal of maintaining a model-agnostic design.
@@ 4.1.3. Baselines
    ENMF (Chen et al., 2020a) introduces mathematical optimization that enable efficient full-data MF without negative sampling.
@@
Both tools and recommendation baselines are pretrained on the first (nk)(n-k) items of each user’s interaction sequence $\mathbf{s}_{[:-k]}$, following the standard training protocol of sequential recommendation for fair performance comparison. Following the previous works (Zhang et al., 2024b; Xu et al., 2025), we also compare AgentDR with language-based approaches:
@@
In addition to these language-based baselines, we compare AgentDR against recommendation tools integrated with Retrieval-Augmented Generation (RAG). Specifically, the LLM is prompted to generate the top 20 items from the 50 candidates retrieved by each tool, conditioned on the user’s behavior sequence. The same hallucination filtering mechanism from Sec. 3.6 is applied for fair comparison. These baselines are represented as (9)-(11) RAGG/S/M for LightGCN, SASRec and SimpleX, respectively. Since most prior LLM-agent works (Zhang et al., 2024b; Wang et al., 2024; Huang et al., 2025c; Zhang et al., 2024a; Wang et al., 2025a; Xu et al., 2025) are unable to rank all items in datasets, we exclude them from the comparison table instead of reporting trivially low full-ranking performance.
@@ 4.2. Semantic Vicinity Evaluation: VDCG
Specifically, we prompt the LLM to rate how well each item in the recommended list aligns with the ground-truth item in terms of relevance, usefulness, and user interest. The rating scale ranges from 0 to 9, where 0 indicates an entirely unrelated item, and 9 represents a perfect semantic match between the item descriptions. These ratings are treated as relevance scores to compute the DCG (Burges et al., 2005) for the list. For normalization, we divide by a customized ideal DCG, calculated from a synthetic ideal list of scores: $p,p/2,...,p/2^{l}]$, where $p$ is the maximum relevance score minus 1, and $l$ is the length of the list. This ideal score distribution reflects the empirical tendency of the LLM to assign higher scores to only a few semantic-related items, following an approximately exponential decay. VDCG thus captures both the semantic vicinity and ranking quality of the predicted list with respect to the user’s true intent.